I'm supposed to simplify each side of equation then solve it. i can't seem to get a answer that works
-5-4+14=7y-12-6y+1

Answer by checkley75(3666)

-5-4+14=7Y-12-6Y+1
5=Y-11
Y=5+11
Y=16 ANSWER.